OpenStudy (anonymous):

So I'm on my last problem, it's already late but they're giving me until 10 of 2... I need to know the mean median mode and range of 1/5, 1/4, 1/3, 1/2, 3/5, 2/3, 3/4, 4/5. I hate having to ask for other peoples help

OpenStudy (amistre64):

can you put them in order for me?

OpenStudy (anonymous):

they're already in order

OpenStudy (amistre64):

lol...ok ;)

OpenStudy (amistre64):

1/2,(median),3/5 what is halfway between 1/2 and 3/5?

OpenStudy (amistre64):

5/10; 6/10 10/20; 12/20 median = 11/20

OpenStudy (amistre64):

range is 4/5 - 1/5 = 3/5

OpenStudy (amistre64):

no mode; modeless, all numbers are equally represented, none appear more than the other

OpenStudy (amistre64):

1/5, 3/5, 4/5 = 8/5 = 32/20 1/4, 3/4, 2/4 = 6/4 = 30/20 1/3, 2/3, = 3/3 62/20 + 20/20 = 82/20(8) 41/10(8) = 41/80 is the mean if i did it right

OpenStudy (anonymous):

can 41/80 be reduced?

OpenStudy (amistre64):

41 is prime

OpenStudy (amistre64):

its a smidge over 1/2

OpenStudy (anonymous):

okey dokey....thanks so much. You just helped me to NOT fail this semester...great greatfulness is coming from me and going to you...lol